OpenStudy (anonymous):

The planet Jupiter of mass 2x10^27 kg revolves round the sun of mass 2x10^30 kg in a circular orbit of radius 7.8x10^11 m ,Calculate the gravitational force and orbit speed of Jupiter

OpenStudy (anonymous):

i've calculated the gravitational force its coming F=4.38x10^23

OpenStudy (anonymous):

but what about orbit speed of jupiter ??

OpenStudy (anonymous):

oh its circular

OpenStudy (anonymous):

root mew over radius

OpenStudy (anonymous):

\[V \approx \sqrt{\frac{ \mu }{ r }}\]

OpenStudy (anonymous):

Sorry mew is just the gravitational constant 6.67 x10^-11 multiplied by the mass of whatevers curving it

OpenStudy (anonymous):

the answer given in my book is 1.3077x10^4 m/s

OpenStudy (anonymous):

the formula is given wrong answer

OpenStudy (anonymous):

i was wrong, sorry, orbits are elliptical in most cases, you take the formula from the first eqation, equate it with kinetic energy, as Jupiter is a huge classical body, and rearrange for V

OpenStudy (anonymous):

well because Jupiter is non-relativistic

OpenStudy (anonymous):

rr \[1/2mv^2 = Gmm/r^2 \]

OpenStudy (anonymous):

x 2, /m, root

OpenStudy (anonymous):

1/2 mv^2 = 4.38x10^23

OpenStudy (anonymous):

1/2*(2x10^27)*v^2 = 4.38x10^23

OpenStudy (anonymous):

(1x10^27)v^2=4.38x10^23

OpenStudy (anonymous):

v^2=4.38x10^-4

OpenStudy (anonymous):

v=0.02092 m/s

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

the answer is wrong again .... the real one is 1.3077x10^4

OpenStudy (anonymous):

the first formula i gave you is correct

OpenStudy (anonymous):

\[V = \sqrt{\frac{ 6.67 \times 10^{-11} \times 2\times10^{30}) }{ 7.8 \times 10^{11}}\]

OpenStudy (anonymous):

That hasnt converted the LaTex on my screen, do you understand?

OpenStudy (anonymous):

i cannot see what have u've posted

OpenStudy (anonymous):

use the first formula, but the big M is the dominant sphere of influence, namely The Sun.

OpenStudy (anonymous):

mew = G times M

OpenStudy (anonymous):

The gravitation force from the sun to jupiter is the centripetal force.\[4,38 10^{23}=\frac{ m _{jup} }{ R}u ^{2}\]

OpenStudy (anonymous):

oh my god of course

OpenStudy (anonymous):

derivation is right now, thanks